The following table gives information on the amount of sugar (in grams) and the calorie count in one serving of a sample of 13 varieties of Kellogg's cereal.

Sugar (grams)

3

15

12

11

8

6

7

0

6

14

20

1

13

Calories

120

200

130

110

100

80

180

80

120

190

190

110

120

The predictive regression equation of the number of calories on the amount of sugar is y^=88.280+5.020x, where x is amount of sugar (in grams) and y is calories.

Calculate the predicted calorie count for a cereal with 16 grams of sugar per serving.

Round your answer to the nearest integer.

?????????? calories
Answer #1

Solution :

Given that ,

y^ = 88.280 + 5.020 x

We plug x = 16 we get ,

Predicted calorie count = 88.280 + 5.020 * 16 = 169 calories